OpenStudy (anonymous):

Find the limit if it exists. lim x->0 (1-cosx)/sinx

OpenStudy (anonymous):

Did you try LH's rule?

OpenStudy (anonymous):

probably haven't had that yet is my guess multiply by \(\frac{x}{x}\) and you should see two limits you know

OpenStudy (noelgreco):

Multiply numerator and denominator by (1+ cos x). Pythagorean identity on top. Easy from there.

OpenStudy (anonymous):

@NoelGreco that seems to be the cleanest solution. +1.

OpenStudy (anonymous):

I agree, thank you!

OpenStudy (noelgreco):

\[\frac{ (1-\cos ^{2}x) }{ \sin x(1+\cos x) }=\frac{ \sin ^{2}x }{ \sin x(1+\cos x)}=\]

OpenStudy (noelgreco):

=0, or I'm up too late.

OpenStudy (anonymous):

it does, thanks again :)